work done in time t on a body of mass m is accerelated from rest to a speed v in time t1 as a function of time t is given by?

Grade:

1 Answers

Arun
6 years ago
W = F × s
and
F = m × a
so
W = m × a × s
"The body is accelerated from 0 to v in time t1"
so
v = 0 + a × t1
a = v / t1
The distance traveled in a certain time t is W = m × (v / t1) × ( (v × t²) / (2 × t1) )
Rearranging the equation,
W = (m × v² × t²) / (2 × t1²)
Or, W = ( m × v² / 2 ) × ( t / t1 )²